A firm has the following revenue and cost functions.

TR = 60 Q - Q2

TC = 12 Q2 +30 Q + 30

Determine the quantity level at which the firm maximizes its total profit.

(Hint: use marginal revenue = marginal cost rule)


1
Expert's answer
2021-01-29T03:30:05-0500

By the definition of the marginal revenue:


"MR=\dfrac{d}{dQ}(TR)=\dfrac{d}{dQ}(60Q-Q^2)=60-2Q."

By the definition the marginal cost:


"MC=\dfrac{d}{dQ}(TC)=\dfrac{d}{dQ}(12Q^2+30Q+30)=24Q+30."

The firm maximizes its total profit when MR=MC:


"60-2Q=24Q+30,""26Q=30,""Q=1.15"